OpenStudy (anonymous):

find the slope of the line with the equation 5 - 3y =2x

OpenStudy (anonymous):

5 - 3y =2x - 3y =2x -5 3y =-2x +5 y =-2/3x +5/3

OpenStudy (anonymous):

y=mx+b where m is the slope

OpenStudy (anonymous):

-2/3 is the slope

OpenStudy (anonymous):

@DoomDude please stop giving away answers

OpenStudy (anonymous):

ok, i was unaware that he did not want the answer

OpenStudy (anonymous):

@DoomDude this is not answers.com - _ -

OpenStudy (anonymous):

It does help me to check my answer. I very mech appreciate the help.

OpenStudy (anonymous):

No, i simply was "helping" him, this website does not say "DO NOT GIVE PEOPLE ANSWERS TO THEIR QUESTIONS" either...

OpenStudy (anonymous):

especially when I mess up the signs sometimes. Sorry to start trouble !

OpenStudy (anonymous):

oh, no worries

OpenStudy (anonymous):

k. Thanks again from a frustrated mainer

OpenStudy (anonymous):

lol, np